### v1

# Chiral superconductivity from parent Chern band and its non-Abelian generalization

Abstract
We propose a minimal model starting from a parent Chern band with quartic dispersion that can describe the spin-valley polarized electrons in rhombohedral tetra-layer graphene. The interplay between repulsive and attractive interactions on top of that parent Chern band is studied. We conduct standard self-consistent mean-field calculations, and find a rich phase diagram that consists of metal, quantum anomalous Hall crystal, chiral topological superconductor, as well as trivial gapped Bose-Einstein condensate. In particular, there exists a topological phase transition from the chiral superconductor to the Bose-Einstein condensate at zero temperature. Motivated by the recent experimental and theoretical studies of composite Fermi liquid in rhombohedral stacked multi-layer graphene, we further generalize the physical electron model to its composite fermion counterpart based on a field theory analysis. We argue that a chiral spin liquid phase can emerge in the vicinity of the chiral superconductor phase of the composite fermion, i.e., the non-Abelian Moore-Read quantum Hall phase. Our work suggests rhombohedral multi-layer graphene as a potential platform for rich correlated topological phases.
Introduction– Topological superconductor (TSC) is one of the most attractive concepts in modern condensed matter physics. It can host non-Abelian quasiparticles, which are the potential building blocks of quantum computation [1, 2, 3, 4, 5, 6]. In the past decades [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24], experimental efforts have been made in searching for chiral topological superconductors, in which the time reversal symmetry is broken. Recent experimental and theoretical progress in graphene-based two-dimensional materials provide a new platform for unconventional superconductors and correlated insulators [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 28, 37, 38, 39, 40, 41, 42]. While the moiré superlattice often plays an important role in those systems [25, 43, 26, 27, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63], it has been shown that the rhombohedral stacked multi-layer graphene exhibits chiral TSC and gate-tunable Chern bands in the absence of moiré pattern, where TSC and quantum anomalous Hall (QAH) insulator are observed at the same electron density [30, 36, 28, 37, 38, 29].
The intertwinement of topology and correlations in graphene-based two-dimensional materials can lead to rich physics. Motivated by this, we propose a minimal model starting from a parent Chern band with quartic dispersion found in rhombohedral tetra-layer graphene [28]. We consider the interplay between repulsive and attractive interactions among electrons. As a concrete example, we choose the Coulomb interaction as the repulsive interaction and the electron-phonon coupling as the attractive one, two of which are the most commonly discussed interactions in solids. Our self-consistent mean-field calculations show a rich phase diagram consists of metal, QAH crystal, chiral TSC, as well as the trivial gapped Bose-Einstein condensate (BEC). In particular, we show there exists a topological phase transition between TSC and BEC phases at zero temperature. Motivated by the recent experimental and theoretical studies of composite Fermi liquid [36, 31, 64, 34, 33, 32, 35], we further provide a field theory analysis that extend our investigation for physical electrons moving under the Berry curvature to composite fermions coupled to dynamical gauge fields, and argue that a chiral spin liquid (CSL) phase [65, 66, 67, 68, 69, 70, 71, 72, 73] can emerge in the vicinity of the chiral TSC phase of the composite fermion [74], i.e., the non-Abelian Moore-Read quantum Hall phase [75]. Our work shed light on investigating varies correlated topological phases in graphene-based materials.
The model– We consider the following minimal model that captures the spin-valley polarized electrons projected to a single parent band,
$H=H_{0}+H_{\rm int}.$ (1)
Here, $H_{0}=\sum_{\bm{k}}\epsilon({\bm{k}})c^{\dagger}_{\bm{k}}c_{\bm{k}}$ is the kinetic term, where $c_{\bm{k}}^{\dagger}$ is the creation operator of an electron with momentum ${\bm{k}}$ in the parent band. According to the spin-valley polarization, the electron is single-flavored. In rhombohedral tetra-layer graphene, the parent band is believed to be a Chern band with quartic dispersion $\epsilon({\bm{k}})=\alpha|{\bm{k}}|^{4}$ [28], and the Brillouin zone (BZ) is chosen to be hexagonal to respect the $C_{3}$ symmetry. The electron filling is fixed to $\nu=1$. The interaction $H_{\rm int}$ can be divided into the repulsive part and the attractive part. As an example, we consider the repulsive interaction from the electron-electron Coulomb interaction $H_{\text{e-e}}$ and the attractive interaction from the electron-phonon coupling $H_{\text{e-ph}}$ [76], both of which are projected to the parent Chern band,
$H_{\text{e-e}}$ $=\sum_{{\bm{k}}_{1}{\bm{k}}_{2}{\bm{k}}_{3}{\bm{k}}_{4}}\tilde{V}_{{\bm{k}}_{1}{\bm{k}}_{2}{\bm{k}}_{3}{\bm{k}}_{4}}c^{\dagger}_{{\bm{k}}_{1}}c^{\dagger}_{{\bm{k}}_{2}}c_{{\bm{k}}_{3}}c_{{\bm{k}}_{4}},$ (2a)
$H_{\text{e-ph}}$ $=-\sum_{{\bm{k}}_{1}{\bm{k}}_{2}{\bm{k}}_{3}{\bm{k}}_{4}}\tilde{U}_{{\bm{k}}_{1}{\bm{k}}_{2}{\bm{k}}_{3}{\bm{k}}_{4}}c^{\dagger}_{{\bm{k}}_{1}}c^{\dagger}_{{\bm{k}}_{2}}c_{{\bm{k}}_{3}}c_{{\bm{k}}_{4}},$ (2b)
with $\tilde{V}_{{\bm{k}}_{1}{\bm{k}}_{2}{\bm{k}}_{3}{\bm{k}}_{4}}=vV({{\bm{k}}_{1}-{\bm{k}}_{4}}){\mathcal{F}}_{{\bm{k}}_{1},{\bm{k}}_{4}}{\mathcal{F}}_{{\bm{k}}_{2},{\bm{k}}_{3}}\delta_{{\bm{k}}_{1}+{\bm{k}}_{2}-{\bm{k}}_{3}-{\bm{k}}_{4}}$ and $\tilde{U}_{{\bm{k}}_{1}{\bm{k}}_{2}{\bm{k}}_{3}{\bm{k}}_{4}}=uU({{\bm{k}}_{1}-{\bm{k}}_{4}}){\mathcal{F}}_{{\bm{k}}_{1},{\bm{k}}_{4}}{\mathcal{F}}_{{\bm{k}}_{2},{\bm{k}}_{3}}\delta_{{\bm{k}}_{1}+{\bm{k}}_{2}-{\bm{k}}_{3}-{\bm{k}}_{4}}$. Here $u,v$ are positive coupling constants that denote the interaction strengths. The electron-electron interaction is given by the gate-screened Coulomb potential $V({\bm{q}})=\tanh{d|{\bm{q}}|}/(d|{\bm{q}}|)$, while the electron-phonon interaction takes the form of $U({\bm{q}})=q_{0}^{2}/(|{\bm{q}}|^{2}+q_{0}^{2})$. In this work, we fix $d=1.0/|{\bm{K}}|$ around the order of the screening length [31, 76], and $q_{0}=0.3|{\bm{K}}|$ satisfying $q_{0}<|{\bm{K}}|$ [77] [^1] In numerical calculations we find the superconducting phase is relatively inert to $d$, but rather sensitive to $q_{0}$ which is consistent with Ref. [77]., where ${\bm{K}}$ denotes the momentum at the $K$-point of the hexagonal BZ. We then define a energy scale for kinetic term $H_{0}$ as $\epsilon_{K}=\alpha|{\bm{K}}|^{4}$, and evaluate $(u,v)$ in terms of $\epsilon_{K}$. The projected form factor ${\mathcal{F}}$ encodes the quantum geometry of the parent band [76],
${\mathcal{F}}_{{\bm{k}},{\bm{k}}^{\prime}}=\exp\bigg{[}-\frac{{\mathcal{B}}_{0}}{4}(|{\bm{k}}^{\prime}-{\bm{k}}|^{2}+2i{\bm{k}}^{\prime}\times{\bm{k}})\bigg{]}.$ (3)
Here, ${\mathcal{B}}_{0}$ is the Berry curvature which can be taken as fixed parameter. Since the quartic parent Chern band can be approximately viewed as the lowest Landau level, in the following we take a uniform Berry curvature ${\mathcal{B}}_{0}=2\pi/\Omega_{\rm BZ}$ corresponding to Chern number equal to one for the parent band, where $\Omega_{\rm BZ}$ is the area of the first BZ.
Self-consistent calculation– The competition between $H_{\text{e-e}}$ and $H_{\text{e-ph}}$ motivates us to investigate the full phase diagram in terms of $u$ and $v$. To see this, we conduct self-consistent mean-field calculations for the model Hamiltonian Eq. (1). We start from both particle-particle and particle-hole mean-field channels, and seek for self-consistent solutions for different $(u,v)$. For simplicity, we define $S({\bm{q}})=vV({\bm{q}})-uU({\bm{q}})$, such that:
$H_{\rm int}=\sum_{{\bm{k}}_{1}{\bm{k}}_{2}{\bm{k}}_{3}{\bm{k}}_{4}}\tilde{S}_{{\bm{k}}_{1}{\bm{k}}_{2}{\bm{k}}_{3}{\bm{k}}_{4}}c^{\dagger}_{{\bm{k}}_{1}}c^{\dagger}_{{\bm{k}}_{2}}c_{{\bm{k}}_{3}}c_{{\bm{k}}_{4}},$ (4)
where $\tilde{S}_{{\bm{k}}_{1}{\bm{k}}_{2}{\bm{k}}_{3}{\bm{k}}_{4}}=S({\bm{k}}_{1}-{\bm{k}}_{4}){\mathcal{F}}_{{\bm{k}}_{1},{\bm{k}}_{4}}{\mathcal{F}}_{{\bm{k}}_{2},{\bm{k}}_{3}}\delta_{{\bm{k}}_{1}+{\bm{k}}_{2}-{\bm{k}}_{3}-{\bm{k}}_{4}}$.
For the particle-particle channel, we decompose the electron momentum as $({\bm{k}}+{\bm{g}})$, where ${\bm{k}}$ is defined in the first BZ, and ${\bm{g}}_{i}$ labels the reciprocal lattice vectors under $C_{3}$ rotation. Under the Hatree-Fock approximation [76], the self-consistent mean-field interaction Hamiltonian can be divided into the Hartree term with ${\bm{k}}_{1}={\bm{k}}_{4}$ and ${\bm{k}}_{2}={\bm{k}}_{3}$,
$H_{\rm H}=\sum_{\begin{subarray}{c}{\bm{k}}_{1}{\bm{k}}_{2}\\ {\bm{g}}_{1}{\bm{g}}_{2}{\bm{g}}_{3}{\bm{g}}_{4}\end{subarray}}2\tilde{S}_{{\bm{k}}_{1}{\bm{k}}_{2}{\bm{k}}_{3}{\bm{k}}_{4}}^{{\bm{g}}_{1}{\bm{g}}_{2}{\bm{g}}_{3}{\bm{g}}_{4}}P_{{\bm{g}}_{1}{\bm{g}}_{4}}({\bm{k}}_{1})c^{\dagger}_{{\bm{k}}_{2}+{\bm{g}}_{2}}c_{{\bm{k}}_{2}+{\bm{g}}_{3}}+E_{\rm H},$ (5)
and the Fock term with ${\bm{k}}_{1}={\bm{k}}_{3}$ and ${\bm{k}}_{2}={\bm{k}}_{4}$,
$H_{\rm F}=-\sum_{\begin{subarray}{c}{\bm{k}}_{1}{\bm{k}}_{2}\\ {\bm{g}}_{1}{\bm{g}}_{2}{\bm{g}}_{3}{\bm{g}}_{4}\end{subarray}}2\tilde{S}_{{\bm{k}}_{1}{\bm{k}}_{2}{\bm{k}}_{3}{\bm{k}}_{4}}^{{\bm{g}}_{1}{\bm{g}}_{2}{\bm{g}}_{3}{\bm{g}}_{4}}P_{{\bm{g}}_{1}{\bm{g}}_{3}}({\bm{k}}_{1})c^{\dagger}_{{\bm{k}}_{2}+{\bm{g}}_{2}}c_{{\bm{k}}_{2}+{\bm{g}}_{4}}+E_{\rm F}.$ (6)
Here, $E_{\rm H}$ ($E_{\rm F}$) is the constant energy shift from Hartree (Fock) contribution, while $\tilde{S}_{{\bm{k}}_{1}{\bm{k}}_{2}{\bm{k}}_{3}{\bm{k}}_{4}}^{{\bm{g}}_{1}{\bm{g}}_{2}{\bm{g}}_{3}{\bm{g}}_{4}}=S({\bm{k}}_{1}+{\bm{g}}_{1}-{\bm{k}}_{4}-{\bm{g}}_{4}){\mathcal{F}}_{{\bm{k}}_{1}+{\bm{g}}_{1},{\bm{k}}_{2}+{\bm{g}}_{4}}{\mathcal{F}}_{{\bm{k}}_{2}+{\bm{g}}_{2},{\bm{k}}_{1}+{\bm{g}}_{3}}\delta_{{\bm{g}}_{1}+{\bm{g}}_{2}-{\bm{g}}_{3}-{\bm{g}}_{4}}$.
Figure 1: (1) Phase diagram of Hamiltonian Eq. (1) from self-consistent calculations. The gray corner shows the possible metallic phase, to the accuracy of our numerics. The yellow block stands for the QAH crystal phase. The blue region denotes the $(p+ip)$-wave TSC, and the purple block represents the trivial gapped BEC phase. (b-d) Schematic phase diagrams for the composite fermion theory Eq. (9), with (b) small Maxwell coupling $\kappa$, (c) negative composite fermion mass $m$, (d) large Maxwell coupling $\kappa$.
For the particle-hole channel, we assume that the pairing happens between electrons with opposite momenta. In this way, the BdG Hamiltonian reads,
$H_{\rm BdG}=\sum_{\bm{k}}\Psi^{\dagger}_{\bm{k}}\begin{pmatrix}[\epsilon({\bm{k}})-\mu]/2&\Delta({\bm{k}})\\ \Delta^{*}({\bm{k}})&[-\epsilon({\bm{k}})+\mu]/2\end{pmatrix}\Psi_{\bm{k}},$ (7)
where $\Psi_{\bm{k}}=(c_{\bm{k}},c^{\dagger}_{-{\bm{k}}})^{T}$ is the Nambu spinor. Here $\mu$ is the chemical potential determined by the self-consistent equation of electron number $N$ at filling $\nu=1$. When $\mu<0$, the BdG Hamiltonian Eq. 7 describes a BEC phase. The superconducting gap
$\Delta({\bm{k}})=\sum_{{\bm{k}}^{\prime}}\frac{1}{2}[S({\bm{k}}-{\bm{k}}^{\prime}){\mathcal{F}}_{{\bm{k}},{\bm{k}}^{\prime}}^{2}-S({\bm{k}}+{\bm{k}}^{\prime}){\mathcal{F}}_{-{\bm{k}},{\bm{k}}^{\prime}}^{2}\langle c_{-{\bm{k}}^{\prime}}c_{{\bm{k}}^{\prime}}\rangle,$ (8)
respects spatial asymmetry $\Delta(-{\bm{k}})=-\Delta({\bm{k}})$ representing the odd parity self-consistent solution of Eq. (7) for the pairing between single-flavored electrons. We also denote $E({\bm{k}})=\sqrt{[\epsilon({\bm{k}})-\mu]^{2}/4+|\Delta({\bm{k}})|^{2}}$ as the dispersion of the Bogoliubov band.
Phase diagram– With above, we conduct self-consistent calculations of Eq. (1) for the particle-particle and the particle-hole channels numerically. We check the convergence of these two channels for each pair of $(u,v)$, and the phase diagram Fig. [1.(a)] is mapped out in terms of $u$ and $v$ varying from 00 to $\epsilon_{K}$. If both channels converge, we pick the energetic favored one. We find that, when both $u$ and $v$ are small ($u,v<0.05\epsilon_{K}$), the system favors a metallic phase, as shown in the gray region. A strong repulsive Coulomb interaction (i.e., large $v$) can drive the system to an insulating phase, with a finite charge gap $\delta$ open at the first BZ edge, as shown in the yellow region. Within the yellow region, the valence band of $H_{\rm HF}$ has non-trivial Chern number and order parameter $P_{{\bm{g}}_{1}{\bm{g}}_{2}}({\bm{k}})$, indicating a QAH crystal phase. Note that, deep in the QAH crystal region, the band is relative flat and the distribution of Berry curvature is rather homogeneous within the first BZ, as shown in the Fig. [2]. We conjecture that the valence band is very close to the lowest Landau level which can give rise to fractional Chern insulators in the presence of moiré potential at fractional $\nu$ [53, 54, 33, 32, 34, 35, 31].
Figure 2: The ratio between band width $w$ and band gap $\delta$ with respect to the interacting strength $u$ and $v$ within the region of QAH crystal phase. The blue, orange, and green curve is plotted at $u=0.2\epsilon_{K}$, $u=0.6\epsilon_{K}$, and $u=1.0\epsilon_{K}$, respectively. One can see that the band width $w$ is small (less than 0.1) compared with the charge gap $\delta$ for all the $(u,v)$ we consider. The inset shows the distribution of Berry curvature within first BZ for $u=0$, $v=0.2\epsilon_{K}$, whose integral gives Chern number $C=1$.
For the $(u,v)$ in the blue region of Fig. [1.(a)], one can find energetic favored convergent solutions for BdG Hamiltonian Eq. (7). The superconducting gap distribution within the first BZ is shown in Fig. [3], whose real (imaginary) part has the profile of $p_{x}$-wave ($p_{y}$-wave) pairing, indicating a chiral TSC phase with $(p+ip)$-wave pairing symmetry [5]. The physical origin for the chiral TSC can be understood as following. For a trivial parent band, due to the Cooper instability, a finite electron-phonon coupling will lead to $s$-wave superconducting phase. For a parent Chern band, the non-trivial Berry curvature will give rise to $p$-wave superconducting phase. This is in accordance with previous work [77, 79]. The vortices in bulk of this phase can trap non-Abelian Majorana zero modes [5].
Figure 3: Superconducting gap distribution within the first BZ with $u=0.1\epsilon_{K}$, $v=0.05\epsilon_{K}$. The dashed line denotes the fermi surface at $\nu=1$. The left and right panel denotes, respectively, the real and imaginary part of $\Delta({\bm{k}})$, which clearly shows the feature of a $(p+ip)$-wave superconductor.
In the purple region of Fig. [1.(a)] we have $\mu<0$ that stands for the BEC phase. The transition between TSC and BEC phases here is rather interesting. The TSC phase has positive chemical potential $\mu$. Deep in the TSC phase, the electrons are dominantly paired close to the Fermi surface, in accordance with the fact that the Bogoliubov band gap is approximately minimized near the first BZ edge deep in the TSC phase. This can be checked for the inset of $u=0.1\epsilon_{K}$ in Fig. [4] showing the corresponding Bogoliubov band dispersion. The Cooper pair is extended in real space. The increasing of electron-phonon interaction $u$ leads to the drop of $\mu$ as well as the decreasing of the Cooper pair size, such that the electrons deep in the Fermi sea starts to pair. Different from the case of $s$-wave, for $(p+ip)$-wave there exists a critical $u^{*}$ where $\mu=0$ and Bogoliubov band gap closes at the $\Gamma$ point. This can be seen that $E({\bm{k}}=0)=0$ when $\mu=0$ and $\Delta({\bm{k}}=0)=0$ for $(p+ip)$-wave. The Chern number for the Bogoliubov band changes from $1$ in TSC phase to 00 in BEC phase, signaling a topological phase transition from the TSC to the trivial gapped BEC phase. In the BEC phase, the more negative $\mu$ is, the more electrons in the Fermi sea are paired and the size of the Cooper pairs are smaller. The gap of Bogoliubov band is minimal at $\Gamma$ point, as shown in the insets with $u=0.47\epsilon_{K}$ and $u=0.6\epsilon_{K}$ in Fig. [4]. Details of the topological phase transition from TSC to BEC phases with fixed $v=0.05\epsilon_{K}$ and varying $u$ are shown in Fig. [4]. We point out that at finite temperature such a phase transition will change to crossover [80].
Figure 4: Transition between the chiral TSC (blue region) and the trivial BEC phase (purple region), with $v=0.05\epsilon_{K}$ and $u\in[0.1,0.6]\epsilon_{K}$. One can see the chemical potential ($\mu$) drop as the increasing of the electron-phonon coupling strength $(u)$. The bottom insets illustrate the band structure of Bogoliubov quasiparticles. The band gap closes at the critical point $(u^{*}\approx 0.33\epsilon_{K},\mu=0)$, indicating the topological phase transition from chiral TSC to trivial BEC phase.
Generalization to composite fermion theory– The phase diagram Fig. [1.(a)] provides insights to study the potential non-Abelian quantum Hall phase in rhomobohedral multi-layer graphene, in which the composite fermi liquid has been studied both experimentally [36] and theoretically [31]. It is natural to extend our model Eq. (1) with physical electrons moving in the background Berry curvature to that of the composite fermion coupled to the dynamical gauge field. The chiral TSC phase for the composite fermion is proposed to be described by the Moore-Read [75, 1] quantum Hall state, which exhibits non-Abelian topological order. Below we generalize the phase diagram Fig. [1.(a)] into the composite fermion case with the help of field theory. The composite fermions in the presence of pairing can be modeled as [74, 81, 82]
$\mathcal{L}$ $=$ $\bar{\psi}i\gamma^{\mu}(\partial_{\mu}-ia_{\mu})\psi-\frac{i}{8\pi}a\mathrm{d}a+m\bar{\psi}\psi+\frac{1}{2\kappa}(\mathrm{d}a)^{2}$ (9)
$+$ $|D_{2a}\Phi|^{2}+r|\Phi|^{2}+t|\Phi|^{4}-\lambda\left(\Phi^{\dagger}\psi^{\beta}\psi_{\beta}+h.c.\right)$
-- $\frac{i}{2\pi}a\mathrm{d}b-\frac{2i}{4\pi}b\mathrm{d}b-\frac{i}{2\pi}b\mathrm{d}A,$
where $m$ is the Dirac mass of the composite fermion $\psi$, which is charged under the dynamical compact $U(1)$ gauge field $a$, and $\kappa$ is the Maxwell coupling of $a$. Coupling constants $t$ and $\lambda$ are set to be positive. Dynamical compact $U(1)$ gauge field $b$ is coupled to the background gauge field $A$ through a mutual Chern-Simons (CS) term, whose flux corresponds to the current of the physical electron. The Maxwell coupling of $b$ is neglected since it is more irrelevant than the self CS term. The complex boson field $\Phi$ describes the pairing of the composite fermion, which is consequently charge-2 under the gauge field $a$. The $\lambda$-term describes $s$-wave superconductivity for the Dirac composite fermion, which is equivalent to the chiral TSC for the non-relativistic composite fermion [74]. This is due to the intrinsic Berry curvature of the Dirac fermion, with a similar mechanism that Berry curvature can turn the $s$-wave superconductor into a chiral $(p+ip)$-wave one in the physical fermion case [77, 83].
The phase diagram of Eq. (9) can be analyzed at the mean-field level. We first consider the non-superconducting phases. This corresponds to $r>0$, where $\Phi$ is gapped and $\psi$ can be integrated out. The resulting topological part of the Lagragian reads
$\mathcal{L}_{\mathrm{QH}}=-\Theta(-m)\frac{i}{4\pi}a\mathrm{d}a-\frac{i}{2\pi}a\mathrm{d}b-\frac{2i}{4\pi}b\mathrm{d}b-\frac{i}{2\pi}b\mathrm{d}A.$ (10)
The topological data, including ground state degeneracy, Hall conductance, anyon charge and statistics can be extracted from the standard $K$-matrix theory [84]. We find that the Hall conductance is $\sigma_{H}=0$ or $\sigma_{H}=1$ for $m$ positive or negative, corresponding to a trivial or an integer quantum Hall (IQH) phase, respectively. Meanwhile, since the absolute value of the determinant of the $K$-matrix for Eq. (10) equals to one, neither of the two phases are topological ordered due to the trivial ground state degeneracy. Thus the $\sigma_{H}=1$ ($\sigma_{H}=0$) phase is topological equivalent to the QAH crystal (gapped BEC) phase in the phase diagram Fig. [1.(a)]. Though the gapped BEC phase exhibits superconductivity while in the $\sigma_{H}=0$ phase the composite fermion does not pair, these two phases are topological equivalent as discussed in the study of proximity effect [85, 5].
When $r<0$, $\Phi$ is condensed, driving the system into superconducting phase for the composite fermion $\psi$. Correspondingly, the $U(1)$ gauge field $a$ is Higgsed to $\mathbb{Z}_{2}$. The pairing potential ensures $\psi$ to remain gapped regardless the sign of $m$, as long as $|m|<|\lambda\left<\Phi\right>|$. Thus, integrating out $\psi$ naively yields the effective Lagrangian
$\mathcal{L}_{\mathrm{MR}}=\frac{1}{2\kappa}(\mathrm{d}a)^{2}+\frac{2i}{2\pi}a\mathrm{d}\tilde{a}-\frac{i}{2\pi}a\mathrm{d}b-\frac{2i}{4\pi}b\mathrm{d}b-\frac{i}{2\pi}b\mathrm{d}A,$ (11)
where $\tilde{a}$ is the Lagrange multiplier introduced such that $a$ is indeed a $\mathbb{Z}_{2}$ gauge field. The Hall conductance in this phase is $\sigma_{H}=1/2$, and the vortex whose current is coupled to $\tilde{a}$ has charge $Q_{\mathrm{v}}=1/4$ and self-statistical angle $\theta_{\mathrm{v}}=\pi/8$. The quasi hole coupled to gauge field $b$ has half charge $Q_{\mathrm{h}}=1/2$ and semionic self-statistics $\theta_{\mathrm{h}}=\pi/2$. These topological data exactly match those in the non-Abelian Moore-Read quantum Hall state. However, Eq. (11) only captures the Abelian part of the anyon statistics. To incorporate non-Abelian statistics, one need to include Majorana zero modes arising naturally in the chiral superconducting phase of the composite fermion. One way to include Majorana after integrating out $\psi$ is to substitute the $\tilde{a}\mathrm{d}a$ term by $(\tilde{a}+\eta i\mathrm{d}\eta/4)\mathrm{d}a$ [86], where $\eta$ describes the Majorana zero mode trapped at the superconducting vortex core. This gives rise to the desired non-Abelian statistics [86] in the Moore-Read state. The non-Abelian Moore-Read quantum Hall phase is corresponding to $(p+ip)$-wave chiral TSC phase in Fig. [1.(a)]. It is important to note that the Pfaffian and the anti-Pfaffian quantum Hall phases with similar non-Abelian statistics can be also achieved by tuning the pairing symmetry of the composite fermion in Eq. (9[74].
In the composite fermion theory, there is an additional phase without a counterpart in the physical electron model and phase diagram Fig. [1.(a)]. To see this, note that when the composite fermion pairs, the gauge field $a$ is a $\mathbb{Z}_{2}$ gauge field which can be confined by increasing the Maxwell coupling $\kappa$ through a 3D Ising transition [87]. When $a$ is confined, the topological part of the theory reads
$\mathcal{L}_{\mathrm{CSL}}=-\frac{2i}{4\pi}b\mathrm{d}b-\frac{i}{2\pi}b\mathrm{d}A=\frac{1}{2}\frac{i}{4\pi}A\mathrm{d}A,$ (12)
where in passing to the second equality the gapped gauge field $b$ is integrated out, yielding a same Hall conductance $\sigma_{H}=1/2$ as the Moore-Read phase discussed previously. This phase is a CSL phase with an Abelian semion whose current is coupled to $b$. Therefore, we conjecture the existence of the CSL phase in the vicinity of the Moore-Read quantum Hall phase with Hall conductance $\sigma_{H}=1/2$ in multi-layer graphene systems.
Equipped with above analysis, we can map out the phase diagram of the composite fermion theory Eq. (9), as shown in Fig. [1 (b-d)]. Fig. [1.(b)] shows when Maxwell coupling $\kappa$ is small, there will be three phases, IQH, Moore-Read, and trivial gapped phases. This is similar to the result in physical electron model shown in Fig. [1.(a)]. When $\kappa$ is strong, the confinement of the ${\mathbb{Z}}_{2}$ gauge field will drive the Moore-Read phase into the CSL phase, as shown in Fig. [1.(d)]. Fig. [1.(c)] is plotted when the composite fermion mass $m<0$, which explictly shows the 3D Ising transition between the Moore-Read phase and the CSL phase.
Discussion– In this work, we discuss the phase diagram of spin-valley polarized electrons under parent Chern band in the presence of both attractive and repulsive interactions. While our study is focused on the parent band with Chern number equals to one, it’s natural to generalize this set up to the case with higher Chern numbers. We conjecture that more ways of asymmetric pairing, such as $(f+if)$-wave and $(h+ih)$-wave, can emerge. Our field theory analysis predicts the existence of chiral spin liquid in the vicinity of the non-Abelian Moore-Read quantum Hall phase, and it will be an interesting future direction to verify it with state-of-the-art numerical methods.
Acknowledgements– We thank Srinivas Raghu, Jing-Yuan Chen, and Zhaoyu Han for discussions on related topics. Y.-Q. W. acknowledges the support from the JQI postdoctoral fellowship at the University of Maryland. Z.-Q. G. is funded through the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, Materials Sciences and Engineering Division under Contract No. DE-AC02-05-CH11231 (Theory of Materials program KC2301).
Supplementary Material
I Hatree-Fock calculation for insulating phase
For the insulating phase we decompose the electron momentum as $({\bm{k}}+{\bm{g}})$, where ${\bm{k}}$ is defined in the first BZ, and ${\bm{g}}$ labels the reciprocal lattice vectors. Under this notation the Hamiltonian reads
$H_{0}=\sum_{{\bm{k}}\in{\rm BZ}}\sum_{{\bm{g}}}\epsilon({\bm{k}}+{\bm{g}})c^{\dagger}_{{\bm{k}}+{\bm{g}}}c_{{\bm{k}}+{\bm{g}}},$ (S1)
and
$H_{\rm int}=\sum_{{\bm{k}}\in{\rm BZ}}\sum_{{\bm{g}}_{1},{\bm{g}}_{2},{\bm{g}}_{3},{\bm{g}}_{4}}$ $S({\bm{k}}_{1}+{\bm{g}}_{1}-{\bm{k}}_{4}-{\bm{g}}_{4}){\mathcal{F}}_{{\bm{k}}_{1}+{\bm{g}}_{1},{\bm{k}}_{4}+{\bm{g}}_{4}}{\mathcal{F}}_{{\bm{k}}_{2}+{\bm{g}}_{2},{\bm{k}}_{3}+{\bm{g}}_{3}}c^{\dagger}_{{\bm{k}}_{1}+{\bm{g}}_{1}}c^{\dagger}_{{\bm{k}}_{2}+{\bm{g}}_{2}}c_{{\bm{k}}_{3}+{\bm{g}}_{3}}c_{{\bm{k}}_{4}+{\bm{g}}_{4}}$ (S2)
×$\times\delta_{{\bm{k}}_{1}+{\bm{g}}_{1}+{\bm{k}}_{2}+{\bm{g}}_{2}-{\bm{k}}_{3}-{\bm{g}}_{3}-{\bm{k}}_{4}-{\bm{g}}_{4}}.$
Under the Hartree-Fock approximation, the interaction Hamiltonian can be decomposed into the Hartree term with ${\bm{k}}_{1}={\bm{k}}_{4}$ and ${\bm{k}}_{2}={\bm{k}}_{3}$:
$H_{\mathrm{H}}=E_{\mathrm{H}}+\sum_{{\bm{k}}\in{\rm BZ}}\sum_{{\bm{g}}_{1},{\bm{g}}_{2},{\bm{g}}_{3},{\bm{g}}_{4}}$ $2S({\bm{g}}_{1}-{\bm{g}}_{4}){\mathcal{F}}_{{\bm{k}}_{1}+{\bm{g}}_{1},{\bm{k}}_{1}+{\bm{g}}_{4}}{\mathcal{F}}_{{\bm{k}}_{2}+{\bm{g}}_{2},{\bm{k}}_{2}+{\bm{g}}_{3}}P_{{\bm{g}}_{1},{\bm{g}}_{4}}({\bm{k}}_{1})c^{\dagger}_{{\bm{k}}_{2}+{\bm{g}}_{2}}c_{{\bm{k}}_{2}+{\bm{g}}_{3}}$ (S3)
×$\times\delta_{{\bm{g}}_{1}+{\bm{g}}_{2}-{\bm{g}}_{3}-{\bm{g}}_{4}},$
and the Fock term with ${\bm{k}}_{1}={\bm{k}}_{3}$ and ${\bm{k}}_{2}={\bm{k}}_{4}$:
$H_{\mathrm{F}}=E_{\mathrm{F}}-\sum_{{\bm{k}}\in{\rm BZ}}\sum_{{\bm{g}}_{1},{\bm{g}}_{2},{\bm{g}}_{3},{\bm{g}}_{4}}$ $2S({\bm{k}}_{1}+{\bm{g}}_{1}-{\bm{k}}_{2}-{\bm{g}}_{4}){\mathcal{F}}_{{\bm{k}}_{1}+{\bm{g}}_{1},{\bm{k}}_{2}+{\bm{g}}_{4}}{\mathcal{F}}_{{\bm{k}}_{2}+{\bm{g}}_{2},{\bm{k}}_{1}+{\bm{g}}_{3}}P_{{\bm{g}}_{1}{\bm{g}}_{3}}({\bm{k}}_{1})c^{\dagger}_{{\bm{k}}_{2}+{\bm{g}}_{2}}c_{{\bm{k}}_{2}+{\bm{g}}_{4}}$ (S4)
×$\times\delta_{{\bm{g}}_{1}+{\bm{g}}_{2}-{\bm{g}}_{3}-{\bm{g}}_{4}}$
Here
$P_{{\bm{g}}_{i},{\bm{g}}_{j}}({\bm{k}})=\left<c^{\dagger}_{{\bm{k}}+{\bm{g}}_{i}}c_{{\bm{k}}+{\bm{g}}_{j}}\right>$ (S5)
is the mean-field ansatz. The constant energy shift $E_{\mathrm{H/F}}$ reads
$E_{\mathrm{H}}$ $=$ $-\sum_{{\bm{k}}\in{\rm BZ}}\sum_{{\bm{g}}_{1},{\bm{g}}_{2},{\bm{g}}_{3},{\bm{g}}_{4}}S({\bm{g}}_{1}-{\bm{g}}_{4}){\mathcal{F}}_{{\bm{k}}_{1}+{\bm{g}}_{1},{\bm{k}}_{1}+{\bm{g}}_{4}}{\mathcal{F}}_{{\bm{k}}_{2}+{\bm{g}}_{2},{\bm{k}}_{2}+{\bm{g}}_{3}}P_{{\bm{g}}_{1},{\bm{g}}_{4}}({\bm{k}}_{1})P_{{\bm{g}}_{2},{\bm{g}}_{3}}({\bm{k}}_{2})\delta_{{\bm{g}}_{1}+{\bm{g}}_{2}-{\bm{g}}_{3}-{\bm{g}}_{4}},$ (S6)
$E_{\mathrm{F}}$ $=$ $\sum_{{\bm{k}}\in{\rm BZ}}\sum_{{\bm{g}}_{1},{\bm{g}}_{2},{\bm{g}}_{3},{\bm{g}}_{4}}S({\bm{k}}_{1}+{\bm{g}}_{1}-{\bm{k}}_{2}-{\bm{g}}_{4}){\mathcal{F}}_{{\bm{k}}_{1}+{\bm{g}}_{1},{\bm{k}}_{2}+{\bm{g}}_{4}}{\mathcal{F}}_{{\bm{k}}_{2}+{\bm{g}}_{2},{\bm{k}}_{1}+{\bm{g}}_{3}}P_{{\bm{g}}_{1}{\bm{g}}_{3}}({\bm{k}}_{1})P_{{\bm{g}}_{2}{\bm{g}}_{4}}({\bm{k}}_{2})\delta_{{\bm{g}}_{1}+{\bm{g}}_{2}-{\bm{g}}_{3}-{\bm{g}}_{4}}.$ (S7)
In calculation we choose total 7 BZs with the first BZ centered and the other 6 BZs packed around it, and 108 ${\bm{k}}$ points within each BZ. The electron filling is set to be $\nu=1$, which means the total number of electron $N$ is equal to the number of ${\bm{k}}$ points in the first BZ, i.e. $N=108$. The ground state energy of the insulating phase under the Hartree-Fock approximation is
$E_{\mathrm{HF}}=\sum_{{\bm{k}}\in\mathrm{BZ}}\bra{{\bm{k}}}H_{0}+H_{\mathrm{H}}+H_{\mathrm{F}}\ket{{\bm{k}}}=-E_{\mathrm{H}}-E_{\mathrm{F}}+\sum_{{\bm{k}}\in\mathrm{BZ}}\epsilon({\bm{k}})$ (S8)
II Mean-field calculation for superconducting phase
For the superconducting phase we assume the pairing happens between electrons with opposite momenta. The mean-field interaction Hamiltonian reads
$H_{\mathrm{SC}}=\sum_{{\bm{k}},{\bm{k}}^{\prime}}S({\bm{k}}-{\bm{k}}^{\prime})\mathcal{F}_{{\bm{k}},{\bm{k}}^{\prime}}^{2}c^{\dagger}_{{\bm{k}}}c^{\dagger}_{-{\bm{k}}}\left<c_{-{\bm{k}}^{\prime}}c_{{\bm{k}}^{\prime}}\right>+h.c.$ (S9)
Define the superconducting gap
$\Delta({\bm{k}})=\sum_{{\bm{k}}^{\prime}}\frac{1}{2}\left[S({\bm{k}}-{\bm{k}}^{\prime})\mathcal{F}_{{\bm{k}},{\bm{k}}^{\prime}}^{2}-S(-{\bm{k}}-{\bm{k}}^{\prime})\mathcal{F}_{-{\bm{k}},{\bm{k}}^{\prime}}^{2}\right]\left<c_{-{\bm{k}}^{\prime}}c_{{\bm{k}}^{\prime}}\right>,$ (S10)
which respects spatial asymmetry $\Delta(-{\bm{k}})=-\Delta({\bm{k}})$ for the pairing of spinless fermion. Then the BdG Hamiltonian reads
$H_{\mathrm{BdG}}=\sum_{\bm{k}}\Psi_{\bm{k}}^{\dagger}\begin{pmatrix}[\epsilon({\bm{k}})-\mu]/2&\Delta({\bm{k}})\\ \Delta^{*}({\bm{k}})&[-\epsilon({\bm{k}})+\mu]/2\end{pmatrix}\Psi_{\bm{k}},$ (S11)
where $\Psi_{\bm{k}}=(c_{\bm{k}},c_{-\bm{k}}^{\dagger})^{\mathbf{T}}$ is the Nambu spinor, and $\mu$ is the chemical potential. The self-consistent equations for the gap $\Delta({\bm{k}})$ and the chemical potential $\mu$ read
$\Delta({\bm{k}})=-\sum_{{\bm{k}}^{\prime}}\frac{\Delta({\bm{k}}^{\prime})}{4E({\bm{k}}^{\prime})}\left[S({\bm{k}}-{\bm{k}}^{\prime})\mathcal{F}_{{\bm{k}},{\bm{k}}^{\prime}}^{2}-S({\bm{k}}+{\bm{k}}^{\prime})\mathcal{F}_{-{\bm{k}},{\bm{k}}^{\prime}}^{2}\right],$ (S12)
$N=\sum_{\bm{k}}1-\frac{\epsilon({\bm{k}})}{E({\bm{k}})},$ (S13)
where $E({\bm{k}})=\sqrt{[\epsilon({\bm{k}})-\mu]^{2}/4+|\Delta({\bm{k}})|^{2}}$ is the dispersion of the Bogoliubov band, which depends on $\mu$ implicitly. $N=108$ is the total number of electrons introduced in the previous section. The total energy of the superconducting state is
$E_{\mathrm{SC}}=\sum_{\bm{k}}\frac{\epsilon({\bm{k}})-\mu}{2}-E({\bm{k}})+\frac{|\Delta({\bm{k}})|^{2}}{2E({\bm{k}})}.$ (S14)
III $K$-matrix calculation and non-Abelian statistics in field theory
As stated in the main text, in the non-superconducting phases of the composite fermion, the topological part of the Lagragian reads
$\mathcal{L}_{\mathrm{QH}}=-\Theta(-m)\frac{i}{4\pi}a\mathrm{d}a-\frac{i}{2\pi}a\mathrm{d}b-\frac{2i}{4\pi}b\mathrm{d}b-\frac{i}{2\pi}b\mathrm{d}A.$ (S15)
Its corresponding $K$-matrix reads
$K_{\mathrm{QH}}=\begin{pmatrix}\Theta(-m)&1\\ 1&2\end{pmatrix},$ (S16)
which does not host topological order since $|\det K|=1$. The charge vector is $\mathbf{p}=(0,1)^{\mathbf{T}}$, and hence the Hall conductance is $\sigma_{H}=\mathbf{p}^{\mathbf{T}}K_{\mathrm{QH}}^{-1}\mathbf{p}=-\Theta(-m)$. Thus $\sigma_{H}=0$ for $m>0$ and $\sigma_{H}=1$ for $m<0$. A topological phase transition occur at $m=0$, where the composite fermiom is gapless.
In the Moore-Read phase, the effective Lagrangian capturing the Abelian part of the anyon statistics is
$\mathcal{L}_{\mathrm{MR}}=\frac{1}{2\kappa}(\mathrm{d}a)^{2}+\frac{2i}{2\pi}a\mathrm{d}\tilde{a}-\frac{i}{2\pi}a\mathrm{d}b-\frac{2i}{4\pi}b\mathrm{d}b-\frac{i}{2\pi}b\mathrm{d}A.$ (S17)
Its corresponding $K$-matrix is
$K_{\mathrm{MR}}=\begin{pmatrix}0&-2&1\\ -2&0&0\\ 1&0&2\end{pmatrix}$ (S18)
with charge vector $\mathbf{q}=(0,0,1)^{\mathbf{T}}$. The Hall conductance in this phase $\sigma_{H}=\mathbf{q}^{\mathbf{T}}K_{\mathrm{MR}}^{-1}\mathbf{q}=1/2$. The vortex carrying “magnetic charge” under the $\mathbb{Z}_{2}$ gauge field $a$ has its current coupled to $\tilde{a}$. Thus, its $l$-vector is $\mathbf{l}_{\mathrm{v}}=(0,1,0)$. The vortex carries charge $Q_{\mathrm{v}}=\mathbf{l}_{\mathrm{v}}^{\mathbf{T}}K_{\mathrm{MR}}^{-1}\mathbf{q}=1/4$ and self-statistical angle $\theta_{\mathrm{v}}=\pi\mathbf{l}_{\mathrm{v}}^{\mathbf{T}}K_{\mathrm{MR}}^{-1}\mathbf{l}_{\mathrm{v}}=\pi/8$. The quasi hole whose current is coupled to gauge field $c$ has $\mathbf{l}_{\mathrm{h}}=(0,0,1)$. Therefore it has half charge $Q_{\mathrm{h}}=\mathbf{l}_{\mathrm{h}}^{\mathbf{T}}K_{\mathrm{MR}}^{-1}\mathbf{q}=1/2$ and semionic self-statistics $\theta_{\mathrm{h}}=\pi\mathbf{l}_{\mathrm{h}}^{\mathbf{T}}K_{\mathrm{MR}}^{-1}\mathbf{l}_{\mathrm{h}}=\pi/2$. The incorporation of Majorana zero modes reads
$\mathcal{L}_{\mathrm{Majorana}}=\frac{2i}{2\pi}\left(\tilde{a}+\frac{1}{4}\eta i\mathrm{d}\eta\right)\mathrm{d}a-i\tilde{a}_{\mu}j^{\mathrm{v}}_{\mu}-ib_{\mu}j^{\mathrm{h}}_{\mu}-\frac{i}{2\pi}a\mathrm{d}b-\frac{2i}{4\pi}b\mathrm{d}b-\frac{i}{2\pi}b\mathrm{d}A,$ (S19)
where $j^{\mathrm{v}}_{\mu}$ and $j^{\mathrm{h}}_{\mu}$ are the currents of the vortices and the quasi holes, respectively, while the Majorana zero modes are described by $\eta$. Equations of motion for $\tilde{a}$ and $b$ gives rise to
$j^{\mathrm{v}}_{\mu}=\frac{1}{\pi}\epsilon_{\mu\nu\lambda}\partial_{\nu}a_{\lambda},\qquad j^{\mathrm{h}}_{\mu}=-\frac{1}{2\pi}\epsilon_{\mu\nu\lambda}\partial_{\nu}(2b_{\lambda}+a_{\lambda}+A_{\lambda}).$ (S20)
Accordingly, the Majorana theory Eq. (S19) can be formally written as
$\mathcal{L}_{\mathrm{Majorana}}=\frac{1}{4}j^{\mathrm{v}}_{\mu}\eta i\partial_{\mu}\eta+\frac{\pi}{8}j^{\mathrm{v}}\frac{i}{\partial}j^{\mathrm{v}}+\frac{1}{4}iA_{\mu}j^{\mathrm{v}}_{\mu}+\frac{\pi}{2}j^{\mathrm{v}}\frac{i}{\partial}j^{\mathrm{h}}+\frac{\pi}{2}j^{\mathrm{h}}\frac{i}{\partial}j^{\mathrm{h}}+\frac{1}{2}iA_{\mu}j^{\mathrm{h}}_{\mu},$ (S21)
after integrating out all the dynamical gauge fields. Here $1/\partial$ is a shorthand of the current-current interaction mediated by gauge fields. The first term describing vortices trapping Majorana zero modes gives rise to the desired non-Abelian statistics. The charge and the Abelian self-statistical angle of vortices, $Q_{\mathrm{v}}=1/4$ and $\theta_{\mathrm{v}}=\pi/8$ respectively, can also be manifestly read off, and similarly for the quasi holes.